4. Value and Debt Consider the following model: (1 -T) Veidt Industries (VI) has one million shares outstanding with a market price of $13.5 per share. VIs permanent debt is $6.25 million. The company and investors pay tax at the following rates Corporate income: 30% . Personal income: T-10% (equity) ·Ti-25% (interest) Although investors expected VI to maintain its existing debt permanently, management plans to announce that VI will repay all of its debt using the proceeds of an equity issue. Management thinks that unlevering will boost VIs share price.(a) (2 points) What is VIs firm value before the announcement? (Vi -) (b) (3 points) What is VIs firm value and share price immediately after the n nouncement but before the new equity is issued and the debt is repaid? (Hints: Think about the implications of the Efficient Market Hypothesis. Start by calculat- ing Vu. Keep in mind that the amount of debt and the number of shares outstanding has NOT yet changed at this point.) (c) (2 points) Was management correct that repaying the debt (unlevering) would n crease the share price? Explain why or why not. (d) (2 points) How many shares will VI have outstanding after the equity issue is com pleted and the debt is repaid in full? (e) (3 points) The Government does not want the tax system to affect capital structure decisions and decides to use the corporate tax rate as a policy tool. By how much should this rate be changed to achieve the Governments desired result? (Assume the government will leave Te 10%, and Ti-25% unchanged.)

Please step by step and thanks for occupying your time.
Answer #1

a. VL = E + D = 1,000,000 *13.5 + 6,250,000 = 19,750,000

b. Now, with The equation VL = VU +[ 1- (1-Tc)(1-Te)/(1-Ti)]* D

19,750,000 = Vu + [1-(1-0.3)(1-0.1)/(1-0.25)]*6,250,000

19,750,000 = Vu + 1,000,000

Vu = 18,750,000

c. No, the management was not correct. Paying up all the debt will result in reduction in the Value of the firm by 1,000,000. This is because of the tax advantage of debt. The risk of investing in the firm becomes lower and hence the return for shareholders will also be less due to lower risk.

d. Number of shares Outstanding = 6,250,000/13.5 + 1,000,000 = 1,462,963 shares

e. The government should try to make (1-Tc)(1-Te)/(1-Ti) = 1

(1-Tc)*(1-0.1)/(1-0.25) = 1

(1-Tc) *1.2 = 1

1-Tc =1/1.2

Tc = 1-0.83333

Tc = 0.1667 = 16.67%

The government should reduce the corporate tax rate to 16.67%